in the aggregate expenditure model for a closed economy assuming investment, government spending and taxes are exogenous, if the marginal prosperity to consume is 0.8, a simultaneous 50 unit increase in government spending and a 20 unit decrease in investment will change equilibrium income by?


1
Expert's answer
2021-05-10T16:40:59-0400

We are assuming that investment, government spending and taxes are exogenous.

Say, Investment = I, Government Spending = G, Tax = T and Autonomous Consumption = C, equilibrum income= Y


Marginal Propensity to Consume is given as:

"MPC = 0.8"

At equilibrum,

"Y=C+I+G"

"Y = C + MPC(Y-T) + I +G."

"Y = C + 0.8(Y) - 0.8(T) + I + G"

"0.2(Y) = C - 0.8(T) + I + G"

 Now, taking differential on both sides, we get

"0.2.\u2206Y = \u2206C - 0.8(\u2206T) + \u2206I + \u2206G"

Now, Government Spending increases by 50 and Investment decreases by 20.

Tax and C remains constant.

Hence,

"\u2206G = 50, \u2206I = -20, \u2206T = \u2206C\u00b0 = 0"

Hence,

"0.2(\u2206Y) = -20 + 50"

"0.2(\u2206Y) = 30"

"∆Y = \frac{30}{0.2}"

"\u2206Y = 150"

The equilibrum income will change by 150 units.